OpenStudy (anonymous):

find the value of x. rt=4x+10 sw=5x-20

OpenStudy (bahrom7893):

I think i know what you mean now.. try this: y = 4x+10 y = 5x-20

OpenStudy (bahrom7893):

therefore: 4x+10=5x-20 20+10=5x-4x 30=x x=30

OpenStudy (anonymous):

that was correct thank you .!